solve the following by any method
y=2x+4
8x-4y=7
solve by any method
8x-4y=16
2x-4=y

Answer by Madgamer107(9)

The two easiest methods to use are substitution and linear combination (I'd show you graphing too, but I'm new to this site, haha, gonna teach myself how to graph soon enough.) Anyway, for linear combination you have to set up the two equations so that you can cancel x's or y's. Watch...
y = 2x + 4
-4y = -8x + 7
You can multiply the top by 4 to make the x's or y's the same amount, and then add to cancel them out...
4y = 8x + 16
-4y = -8x + 7
There's a problem though. Adding the two equations, you'd wind up with 0 = 25, which isn't true. Since both the x's and y's cancel, the two equations represent a set of parallel lines. They will never meet, thus they will never have a common point of intersection. This one is impossible to solve. For the next set...
8x - 4y = 16
2x - 4 = y
Rearrange...
8x - 4y = 16
2x - y = 4
Multiply the bottom by 4 to get common x's
8x - 4y = 16
8x - 4y = 16
For this question, there are an infinite amount of points of intersection. Examining the two equations after simplification, they are in fact the same line. There are several points that would work for both.
Hope I could help!